# Distributed Sign Momentum with Local Steps for Training Transformers

### v1

@@ 4 Experiments
Figure 2: Final validation loss gaps with AdamW under different communication intervals.
@@
Baselines. We benchmark our empirical performance using AdamW. AdamW is the dominant optimizer for the pre-training of Transformers (Liu et al., 2024), and we use its results under synchronized gradient communication as the baseline. We compare the performance gap, measured in validation losses, of Algorithm 1, and SlowMo when base optimizers in both framework are set as AdamW in that SlowMo with AdamW is among the best performing local updating methods (Sun et al., 2024). We omit the comparison with local AdamW with periodic averaging due to its poor performance.
Implementations. We set the total batch size to 480 for all algorithms and run each for 100k steps. In this setup, we fine-tune the recommended parameters for AdamW (both for the base optimizer AdamW in Algorithm 1 and SlowMo): $\beta_{1}=0.9$, $\beta_{2}=0.95$, and a cosine learning rate schedule (Loshchilov and Hutter, 2016) with a 2k-step warm-up. The peak learning rate (LR) is specified in Table 1, and the final LR is set to 0.05×0.05$\times$0.05 × the peak LR. For the global sign momentum step in Algorithm 1, we apply the recommended Lion parameters: $\beta_{1}=0.95$, $\beta_{2}=0.98$, and $\lambda=0.1$ (Liu et al., 2024). We tune the momentum coefficient and global LR for SlowMo, and the global LR for Algorithm 1. Performance is tested on three different sizes of GPT-2 models across varying numbers of workers.
@@
Improved performance. In Figure 1, we present the validation loss curves during the pre-training of the three GPT-2 models, with the communication interval $\tau$ set to 12. In all three experiments, our method consistently outperforms SlowMo by a significant margin. Notably, despite achieving a 12×12$\times$12 × communication reduction, our approach demonstrates only a small final validation loss gap compared to AdamW under perfect synchronized communication, particularly for GPT-2 Small and GPT-2 Medium.
On the effects of communication interval $\tau$. In Figure 2, we illustrate the performance gaps, quantified by increased final validation losses, for the three model pre-training setups under different communication intervals, $\tau=12,24,36$. The scatter plots show that across all model sizes and communication intervals, our method consistently outperforms SlowMo, highlighting the advantages of sign momentum over standard momentum. Additionally, we observe that as $\tau$ increases, the performance gap between our method and SlowMo narrows. Nevertheless, a 10×10$\times$10 × or 20×20$\times$20 × reduction in communication would still offer substantial savings in practical distributed training scenarios.
@@ 4.1 Ablation studies
We conducted several ablation studies to explore the factors contributing to the effectiveness of sign momentum. In all these studies, we used the AdamW optimizer as the base optimizer, as described in the previous section.
Lookahead and signed Lookahead optimizer. We begin by testing whether momentum and signed momentum improve local updating methods in the case of a single worker. To do this, we configure Algorithm 1 with $n=1$, $\beta_{1}=\beta_{2}=\beta$, and $\lambda=0$, reducing it to the signed Lookahead optimizer. Upon that, by replacing (6) with $\bm{x}_{t+1,0}=\bm{x}_{t,0}-\eta\gamma_{t}{\bm{u}}_{t+1}$, we recover the Lookahead optimizer (Zhang et al., 2019). In Figure 5, we plot the validation loss curves during GPT-2 Medium pre-training with communication interval $\tau=48$. The results show that the Lookahead optimizer consistently outperforms the base optimizer AdamW across various choices of the momentum coefficient $\beta$, highlighting the critical role of momentum in training Transformers. Additionally, as shown in Figure 5, the signed Lookahead optimizer also demonstrates improvements over AdamW, albeit with more fluctuations, likely due to the nature of the signed updates. Through these experiments, we demonstrate that both momentum and signed momentum are effective when $n=1$.
Figure 3: Lookahead optimizer: global LR$=1$, $\tau=48$.
Figure 4: Signed Lookahead optimizer: $\tau=24$.
Figure 5: Signed SlowMo: global LR$=3$, $\tau=12$.
@@
We use signed SlowMo to pre-train the GPT-2 Small model with a communication interval of $\tau=12$. As shown in Figure 5, signed SlowMo achieves better final validation losses compared to SlowMo, showing the efficacy of incorporating sign momentum. However, it still underperforms compared to our proposed method, which we hypothesize is due to the acceleration effects provided by using $\beta_{2}>\beta_{1}$.
Figure 6: Global step that mimics AdamW: $\tau=12.$
Adaptive global update. We further compare the global sign momentum method with a global step that mimics AdamW (see Appendix for a description). Balles and Hennig (2018) show that Adam can be regarded as sign momentum with adaptive learning rate depending on relative variance. Our empirical results, as plotted in Figure 6, show that its performance is only comparable with SlowMo, partially indicating that this adaptivity brings limited benefits when used as the global step of local updating methods. We defer its detailed description of this comparisons to the Appendix.
@@ 5 Conclusion and Discussions
In this paper, we have introduced a new framework for distributed sign momentum with local updates, integrating sign momentum with a communication-efficient local update approach. The framework is flexible, accommodating a variety of off-the-shelf base optimizers. Empirical results from training GPT-2 models of various sizes consistently show improvements over baseline algorithms, validating the effectiveness of our approach. Additionally, a series of ablation studies highlights the impact of sign momentum in different scenarios. We also provide theoretical justification for the framework by utilizing a randomized version of the sign operator and proving an $O(1/\sqrt{T})$ convergence rate when using SGD as the base optimizer. To the best of our knowledge, this framework is a novel contribution to the literature.
@@ B.2 Proof of Theorem 2
When $T\geq 36nL^{2}R^{2}\tau^{3}/\eta^{2}$, we have $\gamma L\tau\leq 1/6$, and thus $\frac{1}{\tau T}\sum_{t=0}^{T-1}\sum_{k=0}^{\tau-1}\mathbb{E}\|\nabla f(\bm{x}_{t,k})\|^{2}=O(1/\sqrt{\tau T})$. ∎

### v2

@@ 4 Experiments
Figure 1: Validation loss curves versus communication rounds for GPT-2 small, medium, and large. Communication interval for our Algorithm 1 and SlowMo are set as $\tau=12$.
Figure 2: Validation loss curves versus computation rounds for GPT-2 small, medium, and large. Communication interval for our Algorithm 1 and SlowMo are set as $\tau=12$.
Baselines. We benchmark the empirical performance of our proposed framework, namely, Algorithm 1, using AdamW as the base optimizer, against the state-of-the-art distributed training method with multiple local steps, i.e., SlowMo (Algorithm 5 in Appendix) that also uses AdamW as the base optimizer. AdamW is the dominant optimizer for the pre-training of Transformers (Liu et al., 2024b), and according to (Sun et al., 2024), SlowMo with AdamW as the base optimizer is among the best performing distributed training methods with multiple local steps. We measure practical performance of Algorithm 1 and SlowMo in terms of validation losses, i.e., token-level log perplexity. In general, under the same compute budget, due to significant communication savings, distributed methods with multiple local steps suffer from performance drop compared to those methods that communicate in every computation round. To demonstrate that our proposed Algorithm 1 suffers from much less performance drop, we also report the validation losses of running standalone AdamW optimizer that communicates (all-reduce mini-batch gradients) in every computation round. We omit the comparison with local AdamW, i.e., averaging models over workers after every $\tau$ AdamW steps at each worker, due to its poor performance.
Alg. Com. red. GPT-2 Small 124M GPT-2 Medium 355M GPT-2 Large 770M
Val. Improv. Val. Improv. Val. Improv.
AdamW N.A. 2.917 2.682 2.552
  SlowMo 12×12$\times$12 × 2.993 2.810 2.727
Algorithm 1 12×12$\times$12 × 2.942 5.23% 2.709 10.63% 2.658 7.14%
  SlowMo 24×24$\times$24 × 2.995 2.800 2.733
Algorithm 1 24×24$\times$24 × 2.957 3.8% 2.725 7.79% 2.864 14.00%
  SlowMo 36×36$\times$36 × 3.000 2.818 2.736
Algorithm 1 36×36$\times$36 × 2.972 2.84% 2.750 7.04% 2.714 2.22%
Table 2: Performance comparison of standalone AdamW (with per iteration communcation), SlowMo + AdamW, and Algorithm 1 + AadmW under communcation intervals $\tau=12,24,36$ for GPT-2 small, medium, and large. Com. red. means communication reduction, and Val. denotes final validation losses, Improv. is the improvement of Algorithm 1 over baseline SlowMo in terms of perplexity, i.e., the natural exponential of validation loss improvement.
Implementations. We set the total batch size to 480 for all algorithms, and run each for 100k steps to ensure the same compute budget, as well as the same total number of tokens. In this setup, we fine-tune the recommended parameters for AdamW (both for the base optimizer AdamW in Algorithm 1 and SlowMo): $\beta_{1}=0.9$, $\beta_{2}=0.95$, and a cosine learning rate schedule (Loshchilov and Hutter, 2016) with a 2k-step warm-up. The peak learning rate (LR) is specified in Table 1, and the final LR is set to 0.05×0.05$\times$0.05 × the peak LR. For the global sign momentum step in Algorithm 1, we apply the recommended Lion parameters: $\beta_{1}=0.95$, $\beta_{2}=0.98$, and $\lambda=0.1$ (Liu et al., 2024b). We tune the momentum coefficient and global LR for SlowMo, and the global LR for Algorithm 1. Performance is tested on three different sizes of GPT-2 models across varying numbers of workers.
@@
Performance improvements. In Figure 1 and 2, we present the validation loss curves against communication rounds and computation rounds, respectively, during the pre-training of three GPT-2 models. The communication interval $\tau$ for both our proposed Algorithm 1 and SlowMo is set to 12. Figure 1 and 2 show that our proposed Algorithm 1 saves 12×12$\times$12 × communications compared to standalone AdamW that communicates in every computation round, and achieves comparable performance. In all three experiments, our method consistently outperforms SlowMo by a significant margin, and suffers less than half performance drop compared to the standalone AdamW, particularly for GPT-2 Small and GPT-2 Medium. For example, for GPT-2 medium, Algorithm 1 achieves final validation loss 2.709, while SlowMo only reaches 2.810. The 0.1 validation loss improvement is significant, since according to scaling laws Liu et al. (2024b); Hoffmann et al. (2022), even an 0.05 validation loss improvement can take 2×2$\times$2 × compute to achieve.
On the effects of communication interval $\tau$. In Table 2, we report the final validation losses of training three sized GPT-2 models with various algorithm setups. It shows that Algorithm 1 achieves consistent and robust performance improvements versus SlowMo under different communication intervals, $\tau=12,24,36$, across all model sizes, highlighting the advantages of sign momentum over standard momentum. Additionally, we observe that as $\tau$ increases, the performance gap between our method and SlowMo narrows. Nevertheless, a 10×10$\times$10 × or 20×20$\times$20 × reduction in communication would still offer substantial communication savings in practical distributed training scenarios where per-iteration communication is prohibitively expensive.
More base optimizers. We also test our proposed framework Algorithm 1 using base optimizer Sophia (Liu et al., 2024b), a more recent optimizer shown to outperform AdamW. We train a GPT-2 small model over 4 workers, and compare with SlowMo that also uses Sophia base optimizer under communication interval $\tau=12$, and standalone Sophia with per-iteration mini-batch gradient all-reduce. We use the suggested hyper parameters of Sophia, except tuning its peak LR and accordingly the minimum LR. Similar to previous experiments, we tune momentum coefficient and global LR for SlowMo and global LR for Algorithm 1. We report validation losses in Table 3. Although both Algorithm 1 and SlowMo achieve worse losses than the case of using AdamW as the base optimizer, Algorithm 1 still improves over SlowMo over 5%.
Alg. Com. red. Val. Improv.
Sophia N.A. 2.873
SlowMo 12×12$\times$12 × 3.016
Algorithm 1 12×12$\times$12 × 2.965 5.23%
Table 3: Performance comparison of SlowMo and Algorithm 1 with Sophia as the base optimizer.
@@ 4.1 Ablation studies
We conduct several ablation studies to explore the factors contributing to the effectiveness of sign momentum. In all these studies, we used the AdamW optimizer as the base optimizer, as described in the previous section.
Lookahead and signed Lookahead optimizer. We begin by testing whether momentum and signed momentum improve methods with multiple local steps in the case of a single worker. To do this, we configure Algorithm 1 with $n=1$, $\beta_{1}=\beta_{2}=\beta$, and $\lambda=0$, reducing it to the signed Lookahead optimizer. Upon that, by replacing (7) with $\bm{x}_{t+1,0}=\bm{x}_{t,0}-\eta\gamma_{t}{\bm{u}}_{t+1}$, we recover the Lookahead optimizer (Zhang et al., 2019). We test the Lookahead optimizer using base optimizer AdamW in training GPT-2 medium, under $\tau=48$, global LR = 1, with validation losses reported in Table 4. We also evaluate the signed Loookahead optimizer with AdamW as the base optimizer in training a GPT-2 small model with $\tau=24$, global LR = 6, and the validations losses are reported in Table 5. These results show that both Lookahead and signed Lookahead are effective in improving the performance of base optimizer when $n=1$, highlighting the critical role of momentum in training Transformers.
Alg. $\beta$ Val. Improv.
AadmW N.A. 2.682
Lookahead $0.1$ 2.665 1.71%
Lookahead $0.2$ 2.660 2.22%
Table 4: Performance comparison of AdamW and Lookahead with AadmW as the base optimizer.
Alg. $\beta$ Val. Improv.
AadmW N.A. 2.917
Signed Lookahead $0.6$ 2.896 2.12%
Signed Lookahead $0.8$ 2.890 2.74%
Table 5: Performance comparison of AdamW and signed Lookahead with AadmW as the base optimizer.
@@
We use signed SlowMo to pre-train the GPT-2 Small model with a communication interval of $\tau=12$, and report final validation losses in Table 6. It shows that signed SlowMo achieves better final validation losses compared to SlowMo, showing the efficacy of incorporating sign momentum. However, it still underperforms compared to our proposed Algorithm 1 (best loss 2.942), which we hypothesize is due to the acceleration effects provided by using $\beta_{2}>\beta_{1}$.
Alg. $\beta$ Val. Improv.
AadmW N.A. 2.917
SlowMo 2.993
Signed SlowMo 0.5 2.974 1.92%
Signed SlowMo 0.8 2.970 2.32%
Global AdamW N.A 2.992 0.10%
Table 6: Performance comparison of AdamW; SlowMo, signed SlowMo, and global AdamW with AadmW as the base optimizer.
Adaptive global update. We further compare the global sign momentum method with a global step that mimics AdamW (see Appendix for a description). Balles and Hennig (2018) show that Adam can be regarded as sign momentum with adaptive learning rate depending on relative variance. Its performance, as reported in Table 6, is only comparable with SlowMo, partially indicating that this adaptivity brings limited benefits when used as the global step of local step methods. We defer its detailed description of this comparisons to the Appendix.
@@ 5 Conclusion and Discussions
In this paper, we have introduced a new framework for distributed sign momentum with local steps, integrating sign momentum with a communication-efficient local step approach. The framework is flexible, accommodating a variety of off-the-shelf base optimizers. Comprehensive empirical results from pre-training GPT-2 models of various sizes from scratch consistently show improvements over the baseline method, validating the effectiveness of our approach in scenarios where communicating at every step is impractical. Additionally, a series of ablation studies highlights the impact of sign momentum in different scenarios. We have provided a general theoretical analysis for generic base optimizers, by utilizing a randomized version of the sign operator, and provided an $O(1/\sqrt{T})$ convergence rate when using SGD as the base optimizer. For SGD as base optimizer in particular, under actual sign operator, we have shown an optimal $O(1/T^{1/4})$ convergence that matches the rate of centralized method. To the best of our knowledge, this framework is a novel contribution to the literature.
@@ B.2 Proof of Theorem 2
When $T\geq 36nL^{2}R^{2}\tau^{3}/\eta^{2}$, we have $\gamma L\tau\leq 1/6$, and thus $\frac{1}{\tau T}\sum_{t=0}^{T-1}\sum_{k=0}^{\tau-1}\mathbb{E}\|\nabla f(\bm{x}_{t,k})\|^{2}=O(1/\sqrt{T})$. ∎
B.3 Proof of Theorem 3
B.3.1 Preliminaries
Since we us SGD as the base optimizer, $\bm{d}_{t,k}^{(i)}=\nabla f_{i}(\bm{x}_{t,k}^{(i)},\bm{\xi}_{t,k}^{(i)})$. We define the following notations:
~$\tilde{\nabla}g(\bm{x}_{t,0})$ $=\frac{1}{n}\sum_{i=1}^{n}\sum_{k=0}^{\tau-1}\nabla f_{i}(\bm{x}_{t,k}^{(i)},\bm{\xi}_{t,k}^{(i)}),$ (34)
~$\tilde{\nabla}f(\bm{x}_{t,0})$ $=\frac{1}{n}\sum_{i=1}^{n}\sum_{k=0}^{\tau-1}\nabla f_{i}(\bm{x}_{t,k}^{(i)}).$ (35)
We define the accumulative stochastic gradient noise at $t$-th outer iteration as:
$\bm{\delta}_{t}$ =~$=\tilde{\nabla}g(\bm{x}_{t,0})-\tilde{\nabla}f(\bm{x}_{t,0}).$
We use
$\bm{\epsilon}_{t+1}$ $=\bm{m}_{t+1}-\tau\nabla f(\bm{x}_{t,0}),$
to denote the gradient estimation error. Then, with $\gamma_{t}=\gamma$ for all $t$ an $\beta_{1}=\beta_{2}=\beta$, one has the following outer iteration for $t=0,\ldots,T-1$:
$\bm{m}_{t+1}$ $=\beta\bm{m}_{t}+(1-\beta)\tilde{\nabla}g(\bm{x}_{t,0}),$ (36)
$\bm{x}_{t+1,0}$ $=\bm{x}_{t,0}-\eta\gamma\mathrm{sign}(\bm{m}_{t+1}).$ (37)
We initialize
$\bm{m}_{0}=\mathbf{0}_{d},$
$\bm{x}_{0,0}=\bm{x}_{-1,0}-\eta\gamma\mathrm{sign}(\bm{m}_{0}).$
Hence $\bm{x}_{0,0}=\bm{x}_{-1,0}$ and $\bm{\epsilon}_{0}=-\tau\nabla f(\bm{x}_{t,0})$.
B.3.2 Descent Lemma
Lemma 2.
Under Assumption 1. For the iterates $\{\bm{x}_{t,0}\}$ in Algorithm 1 with $\beta_{1}=\beta_{2}=\beta,\lambda=0$, it satisfies that
$\frac{1}{T}\sum_{t=0}^{T-1}\mathbb{E}\big{[}\|\nabla f(\bm{x}_{t,0})\ \|_{1}\big{]}\leq\frac{f(\bm{x}_{0,0})-f_{*}}{\eta\gamma T}+\frac{2\sqrt{d}}{\tau T}\sum_{t=0}^{T-1}\mathbb{E}\big{[}\|\bm{\epsilon}_{t+1}\|\big{]}+\frac{\eta\gamma dL}{2}.$ (38)
Proof.
The proof of this Lemma is adapted from Sun et al. (2023). From Assumption 1 we have
$f(\bm{x}_{t+1,0})$ $\leq f(\bm{x}_{t,0})+\nabla f(\bm{x}_{t,0})^{\top}(\bm{x}_{t+1,0}-\bm{x}_{t,0})+\frac{L}{2}\|\bm{x}_{t+1,0}-\bm{x}_{t,0}\|^{2}$
(37)${\leq}f(\bm{x}_{t,0})-\eta\gamma\nabla f(\bm{x}_{t,0})^{\top}\mathrm{sign}(\bm{m}_{t+1})+\frac{\eta^{2}\gamma^{2}dL}{2}$
$\leq f(\bm{x}_{t,0})-\eta\gamma\nabla f(\bm{x}_{t,0})^{\top}\mathrm{sign}(\nabla f(\bm{x}_{t,0}))+\eta\gamma\nabla f(\bm{x}_{t,0})^{\top}(\mathrm{sign}(\nabla f(\bm{x}_{t,0})-\mathrm{sign}(\bm{m}_{t+1}))+\frac{\eta^{2}\gamma^{2}dL}{2}$
$=f(\bm{x}_{t,0})-\eta\gamma\|\nabla f(\bm{x}_{t,0})\|_{1}+\eta\gamma\sum_{r=1}^{d}[\nabla f(\bm{x}_{t,0})]_{r}\cdot[\mathrm{sign}([\nabla f(\bm{x}_{t,0})]_{r})-\mathrm{sign}([\bm{m}_{t+1}]_{r})]+\frac{\eta^{2}\gamma^{2}dL}{2}.$
Recall that $\bm{\epsilon}_{t+1}=\bm{m}_{t+1}-\tau\nabla f(\bm{x}_{t,0})$. If $\mathrm{sign}([\nabla f(\bm{x}_{t,0})]_{r})\neq\mathrm{sign}([\bm{m}_{t+1}]_{r})$, then
$|[\bm{\epsilon}_{t+1}]_{r}|=|[\bm{m}_{t+1}]_{r}-\tau[\nabla f(\bm{x}_{t})]_{r}|\geq\tau|[\nabla f(\bm{x}_{t,0})]_{r}|,$
and thus
$[\nabla f(\bm{x}_{t,0})]_{r}\cdot[\mathrm{sign}([\nabla f(\bm{x}_{t,0})]_{r})-\mathrm{sign}([\bm{m}_{t+1}]_{r})]\leq\frac{2|[\bm{\epsilon}_{t+1}]_{r}|}{\tau}.$
It follows that
$f(\bm{x}_{t+1,0})\leq f(\bm{x}_{t,0})-\eta\gamma\|\nabla f(\bm{x}_{t,0})\|_{1}+\frac{2\eta\gamma\sqrt{d}\|\bm{\epsilon}_{t+1}\|}{\tau}+\frac{\eta^{2}\gamma^{2}dL}{2}.$
Summing over $t=0,\ldots,T-1$ and taking expectation, we obtain the desired relation in (38). ∎
B.3.3 Bounding estimation errors
We next try to bound the estimation error $\mathbb{E}[\|\bm{\epsilon}_{t+1}\|]$. By the recursion of $\bm{m}_{t}$ and the definition of $\bm{\delta}_{t}$, for any $10\leq t\leq T-$ 1,
$\bm{m}_{t+1}$ $=\beta\bm{m}_{t}+(1-\beta)\tilde{\nabla}g(\bm{x}_{t,0})$
$=\beta(\bm{\epsilon}_{t}+\tau\nabla f(\bm{x}_{t-1,0}))+(1-\beta)(\bm{\delta}_{t}+\tau\nabla f(\bm{x}_{t,0}))+(1-\beta)(\tilde{\nabla}f(\bm{x}_{t,0})-\tau\nabla f(\bm{x}_{t,0}))$
Then, subtracting $\tau\nabla f(\bm{x}_{t,0})$ from both sides of the above relation gives that
$\bm{\epsilon}_{t+1}$ $=\beta\bm{\epsilon}_{t}+\beta\tau(\nabla f(\bm{x}_{t-1,0})-\nabla f(\bm{x}_{t,0}))+(1-\beta)\bm{\delta}_{t}+(1-\beta)(\tilde{\nabla}f(\bm{x}_{t,0})-\tau\nabla f(\bm{x}_{t,0}))$
$=\beta^{t+1}\bm{\epsilon}_{0}+\tau\sum_{s=0}^{t}\beta^{t+1-s}(\nabla f(\bm{x}_{s-1,0})-\nabla f(\bm{x}_{s,0}))+(1-\beta)\sum_{s=0}^{t}\beta^{t-s}\bm{\delta}_{s}$
$\quad+(1-\beta)\sum_{s=0}^{t}\beta^{t-s}(\tilde{\nabla}f(\bm{x}_{s,0})-\tau\nabla f(\bm{x}_{s,0})).$
Taking Euclidean norm and total expectation on the above relation leads to
$\begin{split}\mathbb{E}[\|\bm{\epsilon}_{t+1}\|]&\leq\beta^{t+1}\mathbb{E}[\|\bm{\epsilon}_{0}\|]+\underbrace{\tau\sum_{s=0}^{t}\beta^{t+1-s}\mathbb{E}\big{[}\|\nabla f(\bm{x}_{s-1,0})-\nabla f(\bm{x}_{s,0})\|\big{]}}_{E_{1}}+\underbrace{(1-\beta)\mathbb{E}\big{[}\|\sum_{s=0}^{t}\beta^{t-s}\bm{\delta}_{s}\|\big{]}}_{E_{2}}\\ &\quad+\underbrace{(1-\beta)\mathbb{E}\big{[}\|\sum_{s=0}^{t}\beta^{t-s}(\tilde{\nabla}f(\bm{x}_{s,0})-\tau\nabla f(\bm{x}_{s,0}))\|\big{]}}_{E_{3}}.\end{split}$ (39)
We next bound $E_{1},E_{2},E_{3}$ one by one.
Bounding $E_{1}$. From $L$-smoothness of $f$, we have, for all $s=0,\ldots,t$,
$\|\nabla f(\bm{x}_{s-1,0})-\nabla f(\bm{x}_{s,0})\|\leq L\|\bm{x}_{s-1,0}-\bm{x}_{s,0}\|\leq\eta\gamma\sqrt{d}L.$
Thus,
$E_{1}\leq\tau\eta\gamma\sqrt{d}L\sum_{s=0}^{t}\beta^{t+1-s}\leq\frac{\tau\eta\gamma\sqrt{d}L}{1-\beta}.$ (40)
Bounding $E_{2}$. By definition,
$E_{2}$ $=(1-\beta)\mathbb{E}\big{[}\|\sum_{s=0}^{t}\beta^{t-s}\bm{\delta}_{s}\|\big{]}$
$\leq(1-\beta)\Big{(}\mathbb{E}[\|\sum_{s=0}^{t}\beta^{t-s}\bm{\delta}_{s}\|^{2}]\Big{)}^{1/2}$
$\leq(1-\beta)\Big{(}\mathbb{E}[\|\sum_{s=0}^{t}\beta^{t-s}(\frac{1}{n}\sum_{i=1}^{n}\sum_{k=0}^{\tau-1}\nabla f_{i}(\bm{x}_{s,k}^{(i)},\bm{\xi}_{s,k}^{(i)})-\frac{1}{n}\sum_{i=1}^{n}\sum_{k=0}^{\tau-1}\nabla f_{i}(\bm{x}_{s,k}^{(i)}))\|^{2}]\Big{)}^{1/2}$
$\leq(1-\beta)\Big{(}\frac{1}{n^{2}}\sum_{s=0}^{t}\beta^{2t-2s}(\sum_{i=1}^{n}\sum_{k=0}^{\tau-1}\mathbb{E}[\|\nabla f_{i}(\bm{x}_{s,k}^{(i)},\bm{\xi}_{s,k}^{(i)})-\nabla f_{i}(\bm{x}_{s,k}^{(i)}))\|^{2}]\Big{)}^{1/2}$ (41)
$\leq\sqrt{\frac{\tau(1-\beta)}{n}}\sigma,$ (42)
where we used the mutual independence of stochastic gradient noise over workers, and the low of total expectations in (41), and used $\sqrt{1-\beta^{2}}>\sqrt{1-\beta}$ in (42).
Bounding $E_{3}$. By Assumption 1 and 3, we have
$\begin{split}E_{3}&=(1-\beta)\mathbb{E}\big{[}\|\sum_{s=0}^{t}\beta^{t-s}\big{[}\frac{1}{n}\sum_{i=1}^{n}\sum_{k=0}^{\tau-1}\big{(}\nabla f_{i}(\bm{x}_{s,k}^{(i)})-\nabla f_{i}(\bm{x}_{s,0})\big{)}\big{]}\|\big{]}\\ &\leq(1-\beta)\frac{L}{n}\sum_{s=0}^{t}\beta^{t-s}\sum_{i=1}^{n}\sum_{k=0}^{\tau-1}\mathbb{E}[\|\bm{x}_{s,k}^{(i)}-\bm{x}_{s,0}\|]\\ &\leq\frac{1}{2}(1-\beta)\tau(\tau-1)\eta LR\sum_{s=0}^{t}\beta^{t-s}\\ &\leq\frac{1}{2}\tau^{2}\eta LR,\end{split}$ (43)
where we used $\|\bm{x}_{s,k}^{(i)}-\bm{x}_{s,0}\|\leq\eta kR$ from the boundedness of $\|\nabla f_{i}(\bm{x}_{s,k}^{(i)},\bm{\xi}_{s,k}^{(i)})\|\leq R$. In light of (40)(42)(43), the relation (39) leads to
$\mathbb{E}[\|\bm{\epsilon}_{t+1}\|]\leq\tau\beta^{t+1}\|\nabla f(\bm{x}_{0,0})\|+\frac{\tau\eta\gamma\sqrt{d}L}{1-\beta}+\sigma\sqrt{\frac{\tau(1-\beta)}{n}}+\frac{1}{2}\tau^{2}\eta LR.$ (44)
B.3.4 Proof of Theorem 3
We are now in a position to show Theorem 3.
Proof.
Using Lemma 2 with bound of $\mathbb{E}[\|\bm{\epsilon}_{t+1}\|]$ in (44), we obtain
$\frac{1}{T}\sum_{t=0}^{T-1}\mathbb{E}\big{[}\|\nabla f(\bm{x}_{t,0})\ \|_{1}\big{]}\leq\frac{f(\bm{x}_{0,0})-f_{*}}{\eta\gamma T}+\frac{2\sqrt{d}}{T(1-\beta)}\|\nabla f(\bm{x}_{0,0})\|+\frac{2dL\eta\gamma}{1-\beta}+2\sigma\sqrt{\frac{(1-\beta)d}{\tau n}}+\eta L\big{(}\tau R\sqrt{d}+\frac{\gamma d}{2}\big{)}.$
Taking $\eta=\frac{1}{LT^{3/4}}$, and $1-\beta=\frac{1}{\sqrt{T}}$, it holds that
$\frac{1}{T}\sum_{t=0}^{T-1}\mathbb{E}\big{[}\|\nabla f(\bm{x}_{t,0})\ \|_{1}\big{]}\leq\frac{L(f(\bm{x}_{0,0})-f_{*})}{\gamma T^{1/4}}+\frac{2\sqrt{d}}{T^{1/2}}\|\nabla f(\bm{x}_{0,0})\|+\frac{2d\gamma}{T^{1/4}}+\frac{2\sigma}{T^{1/4}}\sqrt{\frac{d}{\tau n}}+\frac{\sqrt{d}\tau R+\gamma d/2}{T^{3/4}}.$